The length of a rectangle is 4 in. greater than the width. The perimeter of the rectangle is 24 in. Find the dimensions of the rectangle.

Answer by checkley71(8403)

LENGTH=W+4
WIDTH=W
PERIMETER=2L+2W
24=2(W+4)+2W
24=2W+8+2W
4W=24-8
4W=16
W=16/4
W=4 ANSWER.
L=4+4=8 ANSWER.
PROOF
24=2*4+2*8
24=8+16
24=24